### v3

# Probing LLM Hallucination from Within: Perturbation-Driven Approach via Internal Knowledge

Abstract
LLM hallucination, where unfaithful text is generated, presents a critical challenge for LLMs’ practical applications. Current detection methods often resort to external knowledge, LLM fine-tuning, or supervised training with large hallucination-labeled datasets. Moreover, these approaches do not distinguish between different types of hallucinations, which is crucial for enhancing detection performance. To address such limitations, we introduce hallucination probing, a new task that classifies LLM-generated text into three categories: aligned, misaligned, and fabricated. Driven by our novel discovery that perturbing key entities in prompts affects LLM’s generation of these three types of text differently, we propose SHINE, a novel hallucination probing method that does not require external knowledge, supervised training, or LLM fine-tuning. SHINE is effective in hallucination probing across three modern LLMs, and achieves state-of-the-art performance in hallucination detection, outperforming seven competing methods across four datasets and four LLMs, underscoring the importance of probing for accurate detection.
Probing LLM Hallucination from Within:
Perturbation-Driven Approach via Internal Knowledge
Seongmin Leethanks: Work done during internship at JPMorganChase.12, Hsiang Hsu2, Chun-Fu (Richard) Chen2, Duen Horng (Polo) Chau1 1Georgia Institute of Technology,  2JPMorganChase Global Technology Applied Research seongmin@gatech.edu, {hsiang.hsu,richard.cf.chen}@jpmchase.com, polo@gatech.edu
Figure 1: We introduce a novel hallucination probing task that distinguishes hallucination into misaligned and fabricated text. SHINE, our novel hallucination probing method, achieves best performance in hallucination detection, outperforming 7 methods across 4 datasets and 4 LLMs (Table 2, 3).
1 Introduction
Large language models (LLMs) excel at text generation Wu et al. (2023); Thirunavukarasu et al. (2023) but often produce hallucinations — incorrect or unverifiable content — posing significant risks for practical use Bohannon (2023). Detecting such hallucinations is crucial yet challenging due to their plausible appearance Ji et al. (2023a); Xu et al. (2024b) and diverse types Zhang et al. (2023c); Ji et al. (2023a). One common detection approach involves comparing outputs with external sources (e.g., Wikipedia) Lin et al. (2021); Min et al. (2023); Tang et al. (2024), but they fail when such sources are unavailable. Alternative approaches include fine-tuning LLMs to reject questions likely to produce hallucinations Zhang et al. (2023a); Xu et al. (2024a); Li et al. (2024) or training classifiers to detect them Azaria and Mitchell (2023); Chen et al. (2023); Su et al. (2024), both requiring expensive supervised training on large labeled datasets. To address these limitations, some recent efforts forgo external knowledge or supervised training. These approaches include checking consistency across multiple generations Manakul et al. (2023), prompting LLMs to assess correctness Friel and Sanyal (2023); Zhang et al. (2024b); Chen and Mueller (2024), or estimating LLM’s uncertainty about its generation Zhang et al. (2023b); Farquhar et al. (2024); Kossen et al. (2024); Chen et al. (2024); Yadkori et al. (2024a).
However, the methods above have a key limitation: they do not assess whether an LLM possesses sufficient knowledge to generate accurate responses — an essential factor in ensuring accurate detection. Specifically, consistency checks cannot identify hallucinations that the model repeatedly generates Slobodkin et al. (2023), while prompting approaches may result in random guesses when the LLM lacks knowledge. Moreover, uncertainty-based methods assume all hallucinations stem from uncertainty, neglecting factors such as randomness in token sampling or beam search variability Stahlberg and Byrne (2019); Holtzman et al. (2019). These limitations highlight the critical need to assess the LLM’s knowledge and its alignment with generated content. We address this gap with the following key contributions:
  • A novel hallucination probing task that classifies LLM-generated text into three categories — aligned (LLM has sufficient knowledge and responds faithfully), misaligned (LLM has knowledge but generates contradictions), and fabricated (LLM lacks relevant knowledge). Our categorization refines hallucination types beyond existing work Huang et al. (2023a); Zhang et al. (2023c) (§3), significantly boosting detection performance (Fig. 1).
  • Entity Perturbation Impact discovery, our key finding revealing that perturbing key entities in prompts affects LLM outputs differently depending on the type of hallucination present (§3). While prior studies have identified that hallucinations arise from errors while retrieving knowledge about the key entities Ferrando et al. (2024), they did not leverage this for hallucination detection. We systematically perturb key entities to assess their impact on generated tokens, demonstrating how these perturbations can determine whether the generated tokens stem from the LLM’s knowledge and align with it. This provides the foundation for distinguishing between aligned, misaligned, and fabricated responses for hallucination probing.
  • SHINE, a novel hallucination probing method leveraging our Entity Perturbation Impact discovery to predict whether an LLM possesses sufficient knowledge to generate a faithful response and whether the response aligns with the knowledge (§4, Fig. 3). SHINE operates without external knowledge, supervised training, or LLM fine-tuning. SHINE stands for Systematic Hallucination Inspection with Noisy Entity.
  • Extensive experiments demonstrating the superiority of our method on four modern LLMs (LLaMA2-13B-Chat, LLaMA3-8B-Instruct, Mistral-7B-Instruct, Qwen2.5-7B-Instruct) across four datasets for diverse text generation tasks. Our method effectively classifies aligned, misaligned, and fabricated text (§5.1) and outperforms all existing algorithms on hallucination detection (§5.2), underscoring the importance of hallucination probing.
2 Related Work
Hallucination Probing. Previous studies have examined hallucination causes by inspecting training data, algorithms, and inference process (Sec 3 of Huang et al. (2023a)). Key issues during inference include contradictions with the LLM’s knowledge — arising from its tendency to prioritize user preferences over accuracy, randomness in generation, and dependency on earlier tokens — and overconfident fabrication despite lacking knowledge (Sec 4 of Zhang et al. (2023c)). Based on this, we categorize hallucinations into two classes: (1) misaligned, where the LLM has sufficient knowledge but produces contradictions, and (2) fabricated, caused by the LLM’s lack of knowledge. Our work does not aim to attribute hallucination causes to model components Chen et al. (2025).
Hallucination Detection. Some approaches verify LLM-generated text using external sources or human evaluation, which can be unavailable or costly Lin et al. (2021); Min et al. (2023); Tang et al. (2024). As an alternative, researchers trained classifiers to detect hallucinations Chen et al. (2023); Azaria and Mitchell (2023); Su et al. (2024) or fine-tuned LLMs to avoid potentially hallucinated responses Zhang et al. (2023a); Xu et al. (2024a); Li et al. (2024), but these require training on large labeled datasets. Prompting Friel and Sanyal (2023); Li et al. (2023); Zhang et al. (2024b) or consistency checks Manakul et al. (2023); Yadkori et al. (2024a, b); Chen et al. (2024) may fail against confident fabrications Bengio et al. (2015); Iqbal and Qureshi (2022); Zhang et al. (2023b); Kamath et al. (2024). Uncertainty estimation methods Huang et al. (2023b); Quevedo et al. (2024); Zhang et al. (2023b); Farquhar et al. (2024); Hou et al. (2024); Bakman et al. (2024); Yaldiz et al. (2024); Liu et al. (2025) often overlook hallucinations from randomness or token dependencies Stahlberg and Byrne (2019); Holtzman et al. (2019); Zhang et al. (2024a). We propose a new direction to more accurately identify hallucinations without any external knowledge, training, or impractical assumptions.
3 Discovery: Entity Perturbation Impact for Hallucination Probing
Inspired by prior work that explores GPT-2’s knowledge by perturbing entities in templated prompts Meng et al. (2022), we ask: Is LLM-hallucinated text, ungrounded in model knowledge, as sensitive to perturbations as aligned text? This drives our novel Entity Perturbation Impact discovery, forming the basis of SHINE, our hallucination probing method (§4). To explore this, we extend the analysis to modern LLMs, specifically, LLaMA2-13B-Chat-GPTQ Touvron et al. (2023) (henceforth, LLaMA2), and more flexible prompts that ask factual questions about specific entities, which have been the primary focus of recent research Martino et al. (2023); Sun et al. (2023); Ferrando et al. (2024); Huang et al. (2024). We use the NEC benchmark Liu et al. (2024a), which includes prompts about real and fictional entities across six topics; real-entity prompts yield aligned or misaligned responses, whereas fictional ones generate fabricated responses (details in §A.1).
Below, we outline an LLM’s inference process, then introduce our perturbation and evaluation method, which leads to our key discovery that perturbing key entities affects LLM’s generation of aligned, misaligned, and fabricated responses differently, and ultimately gives rise to SHINE (§4).
Figure 2: Entity Perturbation Impact, measured by KL divergence and $\Delta P_{i}$, varies across aligned, misaligned, and fabricated text. (A) Tokens in aligned and misaligned text, strongly linked to the LLM’s knowledge of the entity, have high KL divergence, while most tokens in fabricated text have near-zero values, indicating weak association with the LLM’s knowledge. The mean KL divergence of fabricated text is significantly lower than aligned and misaligned text. (B) While tokens in aligned and fabricated text generally show negative or near-zero $\Delta P_{i}$, misaligned text has tokens with positive $\Delta P_{i}$, particularly those in the contradictory phrase, leading the mean $\Delta P$ of misaligned text to surpass that of aligned and fabricated text.
LLM inference without perturbation. For a prompt $P$ and text $G$, a tokenizer with a token set $\mathcal{T}$ splits $P$ into $M$ tokens $t_{1:M}=[t_{1},\dots,t_{M}]$ and $G$ into $N$ tokens $t_{M+1:M+N}=[t_{M+1},\dots,t_{M+N}]$, where $t_{i}\in\mathcal{T}$. Each token $t_{i}$ is mapped to a $d$-dimensional embedding vector $\mathbf{e}_{i}\in\mathbb{R}^{d}$ by a token embedding map. An LLM $f$ takes the embedding vector sequence $\mathbf{e}_{1:M+N}=[\mathbf{e}_{1},\dots,\mathbf{e}_{M+N}]$ as an input and computes the probability of each token appearing after each token position $i$; i.e., $f(\mathbf{e}_{1:M+N})=\mathbf{P}_{1:M+N}\in\mathbb{R}^{(M+N)\times\lvert\mathcal{T}\rvert},$ where $\mathbf{P}_{i}\in\mathbb{R}^{\lvert\mathcal{T}\rvert}$ and $\mathbf{P}_{i}(t)$ is the probability of the token $t$ to be generated at the position $i+1$. We summarize these notations in App. H.
Entity Perturbation. For each prompt $P$, we manually identify the key entity $S$, tokenized as $t_{1:K}^{S}$. We locate all token positions $I_{S}$ where $S$ appears in $P$ and the LLM’s response $G$, i.e., $I_{S}=\{i|t_{i:i+K-1}=S\}$. To perturb the embeddings of $S$, we add Gaussian noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,{(3\boldsymbol{\sigma}_{0})}^{2})\in\mathbb{R}^{K\times d}$ to all occurrences of $S$, where $\boldsymbol{\sigma}_{0}$ is the standard deviation of token embeddings in the NEC dataset Meng et al. (2022); i.e., $\hat{\mathbf{e}}_{i:i+K-1}={\mathbf{e}}_{i:i+K-1}+\boldsymbol{\epsilon}$ for $i\in I_{S}$, while leaving other token embeddings unchanged. Then, we input $\hat{\mathbf{e}}_{1:{M+N}}$ to the LLM $f$, obtaining a perturbed probability distribution $\hat{\mathbf{P}}_{1:M+N}=f(\hat{\mathbf{e}}_{1:{M+N}})\in\mathbb{R}^{(M+N)\times\lvert\mathcal{T}\rvert}$. For each token position $i$ in $G$, we evaluate the Entity Perturbation Impact using: (1) Kullback-Leibler (KL) divergence between $\mathbf{P}_{i}$ and $\hat{\mathbf{P}}_{i}$ (i.e., $KL(\mathbf{P}_{i}\|\hat{\mathbf{P}}_{i})$) to capture global effects across the entire vocabulary and (2) change in the generation probability of token $t_{i}$ (i.e., $\Delta P_{i}=\hat{\mathbf{P}}_{i}(t_{i})-{\mathbf{P}}_{i}(t_{i})$), focusing on generated tokens. To mitigate the impact of random Gaussian noise, we repeat this process 10 times with different random seeds.
Our approach differs from methods that detect hallucinations via inconsistencies across generations. While their inconsistencies mean that there is a token position $i$ with multiple inconsistent token $t$s with high generation probability $\mathbf{P}_{i}(t)$, we measure the distance between original and perturbed distributions ${\mathbf{P}}_{i}$ and $\hat{\mathbf{P}}_{i}$.
Experiments. From the NEC dataset, we randomly sample two prompts for each topic for both existent and non-existent entities, yielding 12 aligned, 12 misaligned, and 12 fabricated data points. This analysis focuses on a small subset, as it requires manual identification of key entities; results for full dataset using our automated approach are in §5. We also evaluate each text $G$’s Entity Perturbation Impact by taking mean across all tokens in $G$ except those of the perturbed entity, and refer to the results as the KL divergence and $\Delta P$.
Fig. 2 demonstrates the Entity Perturbation Impact of aligned, misaligned, and fabricated text (additional results in App. F). Fabricated text shows a much lower mean KL divergence (0.21) than aligned (0.92) and misaligned (1.36) text (Fig. 2A). For example, in the aligned response to the prompt “What medical condition(s) is Mucinex used to treat?”, tokens for respiratory and sinus, which are relevant to Mucinex, have high KL divergence, whereas fabricated text’s tokens exhibit near-zero KL divergence, indicating their weak association with the LLM’s knowledge. Moreover, misaligned text contains tokens with positive $\Delta P_{i}$, with a mean $\Delta P$ of 0.03 — higher than aligned (-0.16) and fabricated (-0.03) text, where most tokens have negative or near-zero $\Delta P_{i}$ (Fig. 2B). In the misaligned text about Mucinex, tokens in the misaligned phrase fever, and toothaches show positive $\Delta P_{i}$, suggesting $\Delta P_{i}>0$ as a misalignment indicator.
Key Takeaways. Low Entity Perturbation Impact measured by KL divergence strongly indicates fabricated text, as its tokens are weakly tied to key entities. In contrast, a rise in generation probability after perturbation ($\Delta P>0$) signals misalignment. This is because aligned tokens are grounded in model knowledge of the key entities, so perturbation lowers their probability ($\Delta P<0$), while tokens in fabricated text remains unaffected ($\Delta P\approx 0)$. In §4, we introduce SHINE, a hallucination probing method that leverages Entity Perturbation Impact to distinguish between aligned, misaligned, and fabricated text.
Figure 3: SHINE classifies LLM-generated text into aligned, misaligned, and fabricated through two steps: Model Knowledge Test (§4.1) and Alignment Test (§4.2). The Model Knowledge Test identifies whether the LLM has sufficient knowledge to generate a faithful response by perturbing entities in the prompt and evaluating the impact; fabricated text is differentiated at this stage. For the data that the LLM has enough knowledge, we conduct the Alignment Test to examine whether the text aligns with the LLM’s knowledge.
4 SHINE: Hallucination Probing Method
Leveraging our Entity Perturbation Impact discovery (§3), we develop a two-stage workflow for hallucination probing, consisting of Model Knowledge Test and Alignment Test (Fig. 3). The Model Knowledge Test first assesses whether the LLM has enough knowledge to generate faithful response for the prompt and distinguishes fabricated text from other two types (§4.1). Then, the Alignment Test examines whether the generated text aligns with the LLM’s knowledge, classifying it as either aligned or misaligned (§4.2).
4.1 Model Knowledge Test
Model Knowledge Test builds on our discovery that fabricated text exhibits much lower Entity Perturbation Impact (KL divergence) than aligned or misaligned text. To systematically identify key entities, we detect all entities in a prompt and perturb them with varying strength, determined by attention they receive. The Model Knowledge Test involves four steps: identifying entity, determining perturbation strength, perturbing entity embeddings, and measuring perturbation impact.
Step 1. Entity Identification. Using the named entity and noun chunk extraction of SpaCy Honnibal and Montani (2017), Model Knowledge Test extracts entities $S_{1},\dots,S_{L}$ in $P$ (details in App. G).
Step 2. Perturbation Strength Computation. To prioritize influential entities, we apply stronger perturbations to entities with higher attention. To calculate the attention $att_{l}\in\mathbb{R}^{+}$ of each entity $S_{l}$ Tu et al. (2021), we concatenate tokens in $P$ and $G$ into $t_{1:M+N}$, feed the token sequence into the LLM, and sum the attention values received by the tokens of $S_{l}$ while generating $t_{M+1:M+N}$. We then scale $att_{l}$ into $a_{l}\in[0,1]$:
$\vspace{-1pt}a_{l}={\exp{(k_{att}\cdot att_{l})}\over{\exp{(k_{att}\cdot\max{(att_{1},\dots,att_{L})})}}},\vspace{-1pt}$ (1)
where $k_{att}\geq 0$ controls the impact of attention values; $k_{att}=0$ weighs all entities equally, while larger values emphasize highly attended entities.
Since LLMs tend to fabricate for entities with low frequency in the training data Mallen et al. (2023); Kandpal et al. (2023); Min et al. (2023); Ji et al. (2023b), we further adjust perturbation strength using the generation likelihood of each entity $S_{l}$, which serves as a good proxy for frequency Hartmann et al. (2023). Specifically, we use the negative log-likelihood, scaled by token position to consider later tokens’ importance. For an entity $S_{l}$, tokenized as $t_{1:K_{l}}^{S_{l}}$, the perturbation strength $w_{l}\in\mathbb{R}^{+}$ is calculated as
$\vspace{-2pt}w_{l}=a_{l}\cdot\left(1-{1\over K_{l}}\sum_{k=1}^{K_{l}}{\sqrt{k-1}\log{Pr(t_{k}^{S_{l}}|t_{1:k-1}^{S_{l}})}}\right)^{-1}.$ (2)
Step 3. Entity Perturbation. We perturb $L$ entities by adding noise to their token embeddings. For entity $S_{l}=t_{1:K_{l}}^{S_{l}}$, we identify all token positions $I_{S_{l}}$ where it appears in $(P,G)$: $I_{S_{l}}=\{i|t_{i:i+K_{l}-1}=S_{l}\}$. We add Gaussian noise $\boldsymbol{\epsilon}_{l}^{MKT}\sim\mathcal{N}(0,(\boldsymbol{\sigma}_{l}^{MKT})^{2})\in\mathbb{R}^{K_{l}\times d}$ to all $i\in I_{S_{l}}$, i.e., ${\mathbf{e}}^{MKT}_{i:i+K_{l}-1}$ = $\mathbf{e}_{i:i+K_{l}-1}+\boldsymbol{\epsilon}_{l}^{MKT}$, while leaving other tokens unchanged; the standard deviation $\boldsymbol{\sigma}_{l}^{MKT}=w_{l}\cdot\boldsymbol{\sigma}$, where $\boldsymbol{\sigma}$ is a hyperparameter (details in §5). We input the perturbed embeddings ${\mathbf{e}}^{MKT}_{1:M+N}$ to the LLM to compute the perturbed probability distribution ${{\mathbf{P}}^{MKT}}=f({\mathbf{e}}_{1:M+N}^{MKT})\in\mathbb{R}^{(M+N)\times|\mathcal{T}|}$.
Step 4. Model Knowledge Score Evaluation. We compute KL divergence between $\mathbf{P}$ and ${{\mathbf{P}}}^{MKT}$ at each token position $i$ in the generated text $G$ and refer to it as Model Knowledge Score at the position $i$, which is denoted as $MKS_{i}=KL(\mathbf{P}_{i}\|{\mathbf{P}}^{MKT}_{i})$. As the dependency on the previous tokens can reduce the KL divergence values of some tokens (details in App. F), we evaluate $MKS$, the Model Knowledge Score of $G$, by taking the mean of top-$p^{MKS}$ values of $MKS_{i}$ for $i=M+1,\dots,M+N$ ($0<p^{MKS}\leq 1$):
$MKS={{{1\over\lceil p^{MKS}\cdot N\rceil}{\sum_{i=1}^{\lceil p^{MKS}\cdot N\rceil}{MKS_{i}^{sort}}}}},\vspace{-2pt}$ (3)
where $MKS_{1}^{sort}\geq\dots\geq MKS_{N}^{sort}$ is a sorted permutation of $MKS_{M+1},\dots,MKS_{M+N}$. We repeat this 10 times with different random seeds. If the Model Knowledge Score is lower than a threshold, we classify the text as fabricated; otherwise, we proceed to the Alignment Test (§4.2). We determine the threshold $\tau^{MKS}$ using the Kolmogorov-Smirnov (KS) test Massey Jr (1951) on the validation set: $\tau^{MKS}=\operatorname*{arg\,max}_{\tau\in[0,\infty)}{(\mathbf{F}_{f}^{MKS}(\tau)-\mathbf{F}_{a,m}^{MKS}(\tau))}$, where $\mathbf{F}_{f}^{MKS}$ is the cumulative probability of fabricated data’s Model Knowledge Score and $\mathbf{F}_{a,m}^{MKS}$ is that of aligned or misaligned data.
4.2 Alignment Test
After ensuring the LLM has enough knowledge about $(P,G)$ via Model Knowledge Test, we check if the text $G$ aligns with that knowledge. Specifically, we add noise $\boldsymbol{\epsilon}^{AT}_{l}\sim\mathcal{N}(0,(\boldsymbol{\sigma}_{l}^{AT})^{2})$ to the embeddings of each entity $S_{l}$ to compute the perturbed embeddings ${\mathbf{e}}_{1:M+N}^{AT}$; since Model Knowledge Test has already verified that the model is knowledgeable about $S_{l}$, we do not use the likelihood but adjust $\boldsymbol{\sigma}^{AT}_{l}$ only with $a_{l}$ as $\boldsymbol{\sigma}^{AT}_{l}=a_{l}\cdot\boldsymbol{\sigma}$, where $\boldsymbol{\sigma}$ is the same hyperparameter used in Model Knowledge Test. We then compute the perturbed probability distribution ${\mathbf{P}}^{AT}=f({\mathbf{e}}_{1:M+N}^{AT})\in\mathbb{R}^{(M+N)\times\lvert\mathcal{T}\rvert}$ and $\Delta P_{i}={\mathbf{P}}_{i}^{AT}(t_{i})-\mathbf{P}_{i}(t_{i})$ for $i=M+1,\dots,M+N$.
Since some tokens in the misaligned text can still be aligned with model knowledge, while all tokens in the aligned text should be aligned, we evaluate Alignment Score (AS) by taking the mean of top-$p^{AS}$ $\Delta P_{i}$ values ($0<p^{AS}\leq 1$):
$AS={{1\over{\lceil p^{AS}\cdot N\rceil}}{\sum_{i=1}^{\lceil p^{AS}\cdot N\rceil}{\Delta{P}_{i}^{sort}}}},$ (4)
where $\Delta P_{1}^{sort}\geq\dots\geq\Delta P_{N}^{sort}$ is a permutation of $\Delta{P}_{M+1},\dots,\Delta{P}_{M+N}$ sorted in descending order.
While large negative and large positive $AS$ values indicate alignment and misalignment, respectively, $AS\approx 0$ can imply either. To handle near-zero $AS$, we use SelfCheckGPT Manakul et al. (2023), which takes some computational cost but effectively differentiates text about which a model has enough knowledge. We reduce the computational cost of SelfCheckGPT while achieving high accuracy by running it only on data with near-zero $AS$. Such data is identified by setting lower and upper thresholds, $\tau_{l}^{AS}$ and $\tau_{u}^{AS}$, and selecting data within these thresholds; data points with AS lower than $\tau_{l}^{AS}$ and higher than $\tau_{u}^{AS}$ are directly classified as aligned and misaligned, respectively. The thresholds are determined to maximize correct predictions and minimize incorrect predictions on the validation set as
$\tau_{l}^{AS}=\arg\max_{x}{\left({{1+\mathbf{F}_{a}^{AS}(x)}\over{1+\mathbf{F}_{m}^{AS}(x)}/k_{eff}}\right)}$ (5)
$\tau_{u}^{AS}=\arg\max_{x}{\left({{1+(1-\mathbf{F}_{m}^{AS}(x))}\over{1+(1-\mathbf{F}_{a}^{AS}(x))}/k_{eff}}\right)},$ (6)
where $\mathbf{F}_{a}^{AS}$ and $\mathbf{F}_{m}^{AS}$ are the cumulative probabilities of the Alignment Score for aligned and misaligned data, respectively, and $k_{eff}$ is a hyperparameter that controls efficiency; larger $k_{eff}$ reduces penalization for misclassifications, allowing more data to be classified by $AS$; §5.3 further discusses how $k_{eff}$ balances the efficiency and performance.
5 Experiments
This section demonstrates SHINE’s effectiveness in hallucination probing (§5.1) and detection (§5.2), highlighting its superiority over existing algorithms and the importance of probing. Then, we show how our method improves efficiency while maintaining high performance (§5.3).
5.1 Hallucination Probing
We demonstrate SHINE’s effectiveness on three LLMs: LLaMA2-13B-Chat-GPTQ Touvron et al. (2023), LLaMA3-8B-Instruct Dubey et al. (2024), and Mistral-7B-Instruct v0.3 Jiang et al. (2023), henceforth referred to as LLaMA2, LLaMA3, and Mistral. We use en_core_web_sm model for SpaCy, LLM temperature of 1.0, $\boldsymbol{\sigma}=10\boldsymbol{\sigma}_{0}$, where $\boldsymbol{\sigma}_{0}$ is the standard deviation of token embeddings of each dataset, $k_{att}$ of 0.1, and $k_{eff}$ of 0.1.
5.1.1 Dataset
Our evaluation requires datasets of prompts, LLM-generated responses, and labels (aligned, misaligned or fabricated). Since existing datasets only provide binary labels (hallucinated or not), we create two new datasets—NEC and Biography—featuring trinary labels, using existing datasets Liu et al. (2024a); Min et al. (2023). The NEC dataset contains questions on existent and non-existent concepts and LLM-generated responses, with 359 responses in each category (aligned, misaligned, fabricated) for LLaMA2, 359 for LLaMA3, and 476 for Mistral. The Biography dataset contains real and fake biographies, with 109 aligned, 109 misaligned, and 129 fabricated for LLaMA2; 174 aligned, 174 misaligned, and 174 fabricated for LLaMA3; and 128 aligned, 128 misaligned, and 128 fabricated for Mistral. Each dataset is split evenly into validation and test sets within each label. Dataset construction details can be found in App. A.
5.1.2 Results
[…]
Table 1: Confusion matrix showing SHINE’s hallucination probing performance (%) on LLaMA2, LLaMA3, and Mistral across two datasets: NEC / biography.
Actual Aligned Misaligned Fabricated LLaMA2 Predicted Aligned 75.53 / 81.82 13.53 / 5.45 9.89 / 4.62 Misaligned 11.17 / 9.09 73.53 / 87.27 1.65 / 9.23 Fabricated 13.30 / 9.09 12.94 / 7.27 88.46 / 86.15 LLaMA3 Predicted Aligned 72.93 / 83.91 20.77 / 6.90 6.04 / 1.15 Misaligned 14.92 / 8.05 68.31 / 75.86 8.79 / 1.15 Fabricated 12.15 / 8.05 10.93 / 17.24 85.16 / 97.70 Mistral Predicted Aligned 73.44 / 89.06 11.62 / 3.12 12.45 / 1.56 Misaligned 5.81 / 4.69 65.15 / 87.50 4.15 / 3.12 Fabricated 20.75 / 6.25 23.24 / 9.38 83.40 / 95.31
Table 1 shows the confusion matrices of SHINE on the NEC and biography datasets across the LLaMA2, LLaMA3, and Mistral models. Since no existing methods differentiate text into these three types, we cannot directly compare our approach with others. Overall, SHINE effectively differentiates the three types across both datasets and all three LLMs. For LLaMA2, the Model Knowledge Test correctly detects 88.46% of fabricated data points in the NEC dataset and 86.15% in the biography dataset. The Alignment Test further differentiates aligned and misaligned text, correctly classifying 75.53% of aligned and 73.53% of misaligned data points in the NEC dataset and 81.82% of aligned and 87.27% of misaligned data points in the biography dataset. Similarly, for LLaMA3 and Mistral, SHINE consistently classifies most data points correctly, demonstrating its robust effectiveness in hallucination probing.
5.2 Hallucination Detection
Table 2: SHINE (in bold) outperforms all existing approaches in hallucination detection (%), especially for fabricated text, across the NEC / Biography datasets. For a fair comparison, we adapt SHINE’s predictions to binary by combining its misaligned and fabricated classes into the conventional hallucinated class, and mapping its aligned class to the faithful class.
Aligned Misaligned Fabricated Overall Faithful Hallucinated LLaMA2 SHINE (Ours) 83.69 / 90.58 73.40 / 81.82 86.47 / 94.55 91.21 / 95.38 SelfCheckGPT 63.58 / 76.92 86.70 / 81.82 86.47 / 98.18 17.58 / 50.77 BSDetector 76.69 / 73.80 85.64 / 70.91 82.35 / 98.18 62.09 / 52.31 NLL 65.00 / 57.58 82.98 / 89.09 90.59 / 83.64 21.43 / 0.00 Entropy 58.37 / 55.06 67.55 / 98.18 72.94 / 65.45 34.62 / 1.54 Hallucination Score 48.15 / 68.02 93.62 / 89.09 37.65 / 87.27 13.19 / 27.69 Semantic Entropy 67.46 / 78.32 39.89 / 85.45 97.65 / 61.82 64.84 / 87.69 MARS+SE 67.46 / 76.08 39.89 / 81.82 97.65 / 61.82 64.84 / 84.62 LLaMA3 SHINE (Ours) 82.04 / 90.80 72.93 / 80.46 79.23 / 93.10 93.96 / 98.85 SelfCheckGPT 68.14 / 80.84 78.45 / 87.36 75.96 / 91.95 50.00 / 63.22 BSDetector 81.88 / 80.08 85.64 / 86.21 78.69 / 78.16 81.32 / 75.86 NLL 71.41 / 50.19 66.30 / 74.71 74.86 / 70.11 73.08 / 5.75 Entropy 76.57 / 70.11 75.69 / 55.17 69.40 / 87.36 84.62 / 67.82 Hallucination Score 67.39 / 79.69 65.19 / 75.86 69.95 / 90.80 67.03 / 72.41 Semantic Entropy 60.50 / 77.39 71.27 / 73.56 38.25 / 60.92 71.98 / 97.70 MARS+SE 60.07 / 75.86 72.53 / 74.71 35.16 / 55.17 72.53 / 97.70 Mistral SHINE (Ours) 83.13 / 94.79 73.44 / 89.06 88.38 / 96.77 87.55 / 98.44 SelfCheckGPT 71.09 / 69.27 89.63 / 93.75 83.82 / 98.44 39.83 / 15.63 BSDetector 82.85 / 88.54 88.80 / 93.75 85.06 / 98.44 74.69 / 73.44 NLL 74.14 / 65.10 90.04 / 93.75 84.23 / 95.31 48.13 / 6.25 Entropy 65.01 / 71.35 77.18 / 54.69 65.15 / 100.0 52.70 / 59.38 Hallucination Score 63.21 / 67.19 60.58 / 65.62 72.20 / 100.0 56.85 / 35.94 Semantic Entropy 64.32 / 84.90 73.03 / 78.12 35.68 / 84.38 84.23 / 92.19 MARS+SE 59.75 / 83.85 74.27 / 85.94 25.31 / 79.69 79.67 / 85.94
We evaluate SHINE’s effectiveness in conventional hallucination detection by comparing it with two detection methods (SelfCheckGPT Manakul et al. (2023), BSDetector Chen and Mueller (2024)), and five uncertainty-based methods (negative log-likelihood (NLL) Tonolini et al. (2024), entropy Kadavath et al. (2022), Hallucination Score Zhang et al. (2023b), Semantic Entropy Farquhar et al. (2024), MARS Bakman et al. (2024) [^2] Among MARS variants, we use MARS+SE (Semantic Entropy), which exhibits the best performance.). For a fair comparison, we adapt SHINE’s predictions to binary by combining its misaligned and fabricated classes into the hallucinated class, and mapping aligned to faithful. We do not experiment with methods outside our scope, like those requiring external knowledge Lin et al. (2021); Min et al. (2023); Tang et al. (2024), LLM fine-tuning Zhang et al. (2023a); Xu et al. (2024a); Li et al. (2024), supervised training with hallucination-labeled data Chen et al. (2023); Azaria and Mitchell (2023); Su et al. (2024); Yaldiz et al. (2024), or those with different objective of predicting hallucination likelihood from prompts rather than analyzing generated responses Chen et al. (2024); Ji et al. (2024).
5.2.1 Models and Datasets
To assess detection performance across a wider range of models and tasks, we also evaluate Qwen2.5-7B-Instruct Team (2024) (referred to as Qwen) [^3] We could not experiment with Qwen on the NEC and biography datasets since Qwen’s ability to reject the unanswerable questions resulted in only a few fabricated data points., and two benchmark hallucination detection datasets: HaluEval Li et al. (2023) and TriviaQA Joshi et al. (2017), in addition to the NEC and biography datasets (subsubsection 5.1.1). The HaluEval dataset includes 10,000 tuples of a question, relevant context, and faithful and hallucinated responses, resulting in 10,000 faithful and 10,000 hallucinated data points. The TriviaQA dataset contains trivia questions that can be verified using Wikipedia, resulting in 257 faithful and 318 hallucinated data points for LLaMA2, 263 faithful and 318 hallucinated for LLaMA3, 268 faithful and 318 hallucinated for Mistral, and 209 faithful and 318 hallucinated for Qwen. Further details are provided in App. A. We use half of each dataset as validation data, and the other half as test data. As all HaluEval prompts include relevant context, we conduct only the Alignment Test without the Model Knowledge Test. For TriviaQA, which does not distinguish between misaligned and fabricated data and thus cannot decide $\tau^{MKS}$, we set $\tau^{MKS}$ for each model to the mean from the NEC and biography datasets.
5.2.2 Detection Results
Table 2 shows the mean and class-wise accuracies444 We could not compare receiver operating characteristic (ROC) curves as SHINE detects hallucinations using two scores, not one. The ROC curves for the Model Knowledge Test can be found in App. E. of each method on the NEC and biography datasets; class-wise accuracies measure the ratio of aligned text predicted as faithful, misaligned text predicted as hallucinated, and fabricated text predicted as hallucinated; mean accuracy is the average of these three values. SHINE consistently achieves the highest mean accuracies across all models and datasets. Comparing SHINE with the two hallucination detection methods (SelfCheckGPT, BSDetector), which misclassify much of the fabricated text, underscores the importance of differentiating fabricated text from misaligned text for accurate detection. For instance, for LLaMA2, SHINE correctly classifies 91.21% of fabricated text in the NEC dataset, much higher than the 17.58% and 62.09% by SelfCheckGPT and BSDetector, respectively; the underperformance of SelfCheckGPT and BSDetector are attributable to LLM’s consistent generation of fabricated text Slobodkin et al. (2023). The underperformance of uncertainty-based methods (NLL, entropy, Hallucination Score, Semantic Entropy, MARS+SE) aligns with findings that LLMs often exhibit overconfidence in fabricated content Slobodkin et al. (2023). SHINE’s superior performance extends to LLaMA3 and Mistral models, reaffirming its effectiveness.
Table 3 demonstrates SHINE’s superiority on the HaluEval and TriviaQA datasets across four models, showcasing its effectiveness with and without relevant context provided in the prompt. Notably, SHINE is the only method that correctly classifies certain hallucinated data in TriviaQA. For example, for the question “Melanie Molitor is the mom of which tennis world No 1?", LLaMA2 consistently generates the incorrect response “Serena Williams” across ten generations. While all other methods misclassify this as faithful, SHINE correctly classifying it as fabricated (i.e., hallucinated) with a Model Knowledge Score of 0.0369.
Table 3: SHINE (bolded) outperforms all existing methods in hallucination detection (%) on HaluEval/TriviaQA.
Type Overall Faithful Hallucinated LLaMA2 SHINE (Ours) 86.69 / 93.20 83.32 / 96.90 90.96 / 89.51 SelfCheckGPT 76.56 / 92.74 77.30 / 98.45 75.82 / 87.04 BSDetector 80.94 / 88.22 84.96 / 84.62 76.92 / 91.82 NLL 62.65 / 75.29 45.66 / 73.85 79.64 / 76.73 Entropy 63.29 / 72.18 51.50 / 77.69 78.08 / 66.67 Hallucination Score 55.06 / 72.53 11.90 / 81.54 98.22 / 63.52 Semantic Entropy 76.79 / 90.63 76.72 / 86.92 80.96 / 94.34 MARS+SE 76.76 / 90.63 72.82 / 86.92 80.70 / 94.34 LLaMA3 SHINE (Ours) 87.95 / 87.34 81.00 / 87.88 94.90 / 86.79 SelfCheckGPT 78.77 / 87.15 80.08 / 89.39 77.46 / 84.91 BSDetector 84.96 / 86.69 86.56 / 84.09 83.36 / 89.31 NLL 32.00 / 57.58 41.20 / 48.48 22.80 / 66.67 Entropy 82.79 / 63.20 78.86 / 62.88 86.72 / 63.52 Hallucination Score 70.28 / 71.49 68.38 / 56.82 72.18 / 86.16 Semantic Entropy 81.12 / 83.68 81.98 / 81.82 80.26 / 85.53 MARS+SE 81.12 / 84.06 81.98 / 82.58 80.26 / 85.53 Mistral SHINE (Ours) 83.13 / 89.51 89.46 / 86.57 76.80 / 92.45 SelfCheckGPT 77.97 / 89.14 78.88 / 85.82 77.06 / 92.45 BSDetector 81.92 / 85.19 85.32 / 83.58 78.52 / 86.79 NLL 79.05 / 78.25 64.68 / 87.31 93.42 / 69.18 Entropy 67.65 / 72.58 55.08 / 67.16 80.22 / 77.99 Hallucination Score 43.78 / 68.36 75.82 / 45.52 11.74 / 91.19 Semantic Entropy 79.18 / 82.46 78.60 / 76.87 79.76 / 88.05 MARS+SE 79.35 / 82.63 78.98 / 77.21 79.72 / 88.05 Qwen SHINE (Ours) 89.40 / 85.81 82.74 / 92.38 96.06 / 79.25 SelfCheckGPT 76.51 / 85.18 76.60 / 92.38 76.42 / 77.99 BSDetector 85.51 / 84.81 84.22 / 79.05 86.80 / 90.57 NLL 59.47 / 48.60 25.94 / 69.52 93.00 / 27.67 Entropy 64.27 / 50.07 45.52 / 48.57 83.12 / 51.57 Hallucination Score 71.76 / 68.18 55.64 / 53.33 87.88 / 83.02 Semantic Entropy 80.15 / 83.90 78.58 / 86.67 81.72 / 81.13 MARS+SE 80.15 / 83.90 78.58 / 86.67 81.72 / 81.13
5.3 Efficiency of SHINE
SHINE incorporates the hyperparameter $k_{eff}$ to control the efficiency of the Alignment Test by changing the number of data points inspected by SelfCheckGPT. While SelfCheckGPT takes 60.44 seconds to classify a single NEC example using LLaMA2 on an NVIDIA A10G GPU, SHINE’s Alignment Test completes the task in under 2 seconds — over 30$\times$ faster. We evaluate SHINE’s performance and runtime on the NEC dataset and the LLaMA2 model for $k_{eff}$ values ranging from 0.1 to 10 (Fig. 1), with comparisons to other methods. Results for other datasets and models, which follow similar trends, and further analysis are in App. C. SHINE’s performance improves with increased computation time, indicating that SelfCheckGPT often helps the Alignment Test. Even at low $k_{eff}$ values, SHINE outperforms BSDetector, the best-performing compared method, highlighting its superiority in both efficiency and accuracy.
6 Conclusion
We develop SHINE, a novel method to classify LLM-generated text into aligned, misaligned, and fabricated to identify the causes of hallucinations and improve existing detection methods. Model Knowledge Test effectively detects fabricated text and Alignment Test further differentiates aligned and misaligned text. We aim to develop a more efficient and effective technique and to evaluate our method on a broader range of tasks.
7 Limitations
Our hallucination probing task consists of three categories, each of which can be further refined into sub-classes. For example, the ‘misaligned’ class can be divided based on its underlying causes, such has randomness or dependencies on preceding tokens Stahlberg and Byrne (2019); Holtzman et al. (2019); Zhang et al. (2024a). Also, while SHINE opts to focus on hallucinations from the erroneous knowledge retrieval about key entities in prompts — given their prevalence — expanding this scope to other types of hallucinations (e.g., input-conflicting or context-conflicting hallucinationsZhang et al. (2023c)) and prompts (e.g., arithmetic problems, reasoning-based questions) would further enhance its applicability. Additionally, following prior work, we focus on entity keywords, but would like to highlight the potential to extend our approach to non-entity keywords.
8 Potential Risks
We utilize LLMs as proxy knowledge bases, classifying the text that aligns with LLMs’ knowledge as faithful. However, since LLMs may contain outdated or potentially risky information, careful scrutiny would be essential when interpreting their outputs in practical applications, where inaccurate or risky information could have significant consequences.
Acknowledgements
ChatGPT was used to check grammar and spelling of this paper.
Disclaimer
This paper was prepared for informational purposes by the Global Technology Applied Research center of JPMorgan Chase & Co. This paper is not a product of the Research Department of JPMorgan Chase & Co. or its affiliates. Neither JPMorgan Chase & Co. nor any of its affiliates makes any explicit or implied representation or warranty and none of them accept any liability in connection with this paper, including, without limitation, with respect to the completeness, accuracy, or reliability of the information contained herein and the potential legal, compliance, tax, or accounting effects thereof. This document is not intended as investment research or investment advice, or as a recommendation, offer, or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction.
Appendix
Appendix A Dataset Details
In this section, we elaborate on how we construct the datasets used for hallucination probing and detection. All datasets consist of tuples of (prompt, LLM response, label), where the label is one of aligned, misaligned, and fabricated. We assume that all prompts are in English and unambiguous, having no synonyms in the context given in the prompt. We provide example data points in Table 4, Table 5, Table 6, and Table 7.
A.1 NEC dataset
NEC dataset Liu et al. (2024a) consists of 2,073 questions about existent and 2,078 questions about non-existent concepts covering various topics (foods, sports, countries, animals, medicines, generic) curated to examine LLMs’ behaviors asked about unknown questions. While we can ensure that LLM responses for the questions about non-existent concepts are fabricated, we cannot guarantee that the LLM has knowledge to answer all questions about the existent concepts. To identify the questions about which LLM has enough knowledge, we leave only 1,369 questions about the existent concepts on Wikipedia. For each of these questions, we generate 10 succinct responses without reasoning with the studied LLM; we use system instruction “You are a helpful AI assistant that answers #Question#. Keep your answer short and succinct.” and apply the LLM’s chat template if available. We then evaluate the correctness of each response by providing LLaMA3-ChatQA-1.5-8B Liu et al. (2024b), which excels at retrieval-augmented generation, with the question, response, and Wikipedia article about the concept [^5] We use Wikipedia article for only dataset construction. To test our methods, we do not use any external knowledge., following the prompt template adapted from existing literature Manakul et al. (2023); Gu et al. (2024):
Is the answer to the question supported by the article? Answer ‘‘Yes’’ or ‘‘No’’.
Article: {Wikipedia article}
Question: {Question}
Answer: {Response}
If more than 80% (8 out of 10) of the LLM responses are supported by the Wikipedia article, we consider the question to be known and include the question in our dataset; we examine correctness of multiple responses for one question to ensure high quality of the correctness labels, while threshold of 80% accounts for LLMs’ inherent inconsistency due to sampling Saxena et al. (2024). Then, for each known question, we sample one of the supported responses and add the tuple of the (question, response, aligned) in our dataset. To generate misaligned response for each question, we prompt the studied LLM to generate text that contradicts the aligned response and Wikipedia article; we use the same prompt designed to induce factual contradiction to construct HaluEval Li et al. (2023), a hallucination benchmark dataset. As a result, we collect 359 data points for each of the aligned, misaligned and fabricated categories for the LLaMA2 model, 358 data points for each category for LLaMA3, and 476 for each category for Mistral. Within each topic and label, the data is randomly split evenly into validation and test sets [^6] During our experiments, we evaluated SHINE with 5 different validation-test splits, and the variation across different splits is marginal. Therefore, we decided to experiment with one split for each dataset..
A.2 Biography dataset
Biography dataset Min et al. (2023) consists of people’s names on Wikipedia so that we can prompt LLMs to tell a biography of each person. As an LLM’s knowledge about each name greatly varies Min et al. (2023), we identify the names that the LLM knows well by generating a biography, masking out the name from the biography, and asking the LLM to guess what the masked name is; the LLM would be able to correctly recover the name only when the LLM-generated biography contains a lot of information so that it can uniquely indicate the person. However, there is a possibility that the LLM contains a little information about the people that are labeled as unknown. To create fake people about which the LLM would completely fabricate, we assign random jobs which do not match with their true jobs. As a result, we collect LLM-unknown names paired with a wrong job; we pair the LLM-known names with their correct jobs for consistency. From the collected name-job pairs, we create questions of “Tell me a biography of the [job] [name].” and generate responses using the LLM. While we label biography of fake people as fabricated, we take additional care to generate aligned biographies of the LLM-known people as biography is easily hallucinated Min et al. (2023); we collect correct fact atoms that the LLM knows by generating 10 biographies, atomizing each of them, and verifying the correctness of each fact atom by prompting LLaMA3-ChatQA-1.5-8B with the same prompt template used for the NEC dataset. Then, we prompt the studied LLM to generate a biography based on the correct atoms and label it as aligned. To generate misaligned biographies, we change 50% of the fact atoms to be factually contradictory to the original atom and Wikipedia article using the prompting technique from HaluEval Li et al. (2023), and construct biographies based on them. As a result, 109 aligned, 109 misaligned, 129 fabricated biographies are generated for the LLaMA2 model, 174 aligned, 174 misaligned, 174 fabricated biographies for LLaMA3, and 128 aligned, 128 misaligned, 128 fabricated biographies for Mistral; we check whether the generated biographies do not contain any sensitive contents manually and by assessing with LLaMA3-8B-Instruct. We use random half of these data as validation data and the other half as test data.
A.3 HaluEval dataset
HaluEval dataset consists of 10,000 tuples of a question, relevant context, and faithful and hallucinated responses. This results in 10,000 faithful and 10,000 hallucinated data points.
A.4 TriviaQA dataset
TriviaQA contains 318 Wikipedia-verified trivia questions and a set of correct answers for each question. For each model, we generate 10 responses for each question and check the factualness of each response by comparing it with the given set of the correct answers. If more than half of the responses are wrong, we sample one of the wrong responses and label the pair of the question and the sampled response as hallucinated, regarding the model incapable of correctly answering the question. Otherwise, we sample one correct and one wrong responses to generate one faithful and one hallucinated data point for each question; if 10 responses are all correct, we prompt the model to generate a wrong answer following the existing hallucination-inducing strategy Li et al. (2023). As a result, we collect 257 faithful and 318 hallucinated data points for the LLaMA2 model, 263 faithful and 318 hallucinated for LLaMA3, 268 faithful and 318 hallucinated for Mistral, and 209 faithful and 318 hallucinated for Qwen. Half of these data points are randomly sampled as validation data, and the remaining is used as test data.
Table 4: Example data points in the NEC dataset generated for LLaMA2
Prompt LLM-generated text Type
What is the habitat of Pika? Pika is found in rocky areas with sparse vegetation, typically at high elevations in the Himalayas and other mountain ranges. Aligned
What is the habitat of Pika? Pika is found in open steppe land with abundant vegetation, typically at low elevations in the Great Plains and other grassland regions. Misaligned
What is the habitat of Hornoda? Hornoda is found in the deserts of North Africa and the Middle East. Fabricated
What equipment and playing environment are required for Iceclimbing? Ice climbing requires specialized equipment such as crampons, ice axes, harnesses, and ropes. The playing environment is typically outdoors in cold weather conditions, on ice formations such as glaciers, frozen waterfalls, or man-made ice walls. Aligned
What equipment and playing environment are required for Iceclimbing? Ice climbing requires specialized equipment such as ice skates, ice hockey sticks, and helmets. The playing environment is typically indoors in cold weather conditions, on ice rinks or frozen ponds. Misaligned
What equipment and playing environment are required for Cycloling? Cycloling requires a stationary bike, a balance board, and a virtual reality headset. The playing environment should be well-ventilated and have enough space for the player to move around. Fabricated
Table 5: Example data points in the biography dataset generated for LLaMA2
Prompt LLM-generated text Type
Tell me a bio of the film director Ram Gopal Varma. Ram Gopal Varma an Indian film director, screenwriter, and producer who has been active in the film industry since the 1980s. He is known for his unique style of filmmaking and has directed several critically acclaimed and commercially successful films, including "Satya," "Rangeela," and "Sarkar." Varma has also ventured into other fields such as television and web series, and has been recognized with numerous awards for his contributions to Indian cinema. Aligned
Tell me a bio of the film director Ram Gopal Varma. Ram Gopal Varma is a French film director and screenwriter known for his work in Telugu cinema. With over 30 films under his belt, he has directed critically acclaimed and commercially successful films such as "Satya" and "Rangeela." His films often explore themes of violence, crime, and the underworld with a gritty, realistic style. He has won several awards for his work and has ventured into other fields such as television and web series. Ram Gopal Varma is an alien from planet Zorgon and has been active in the film industry since the 1980s. Misaligned
Tell me a bio of the actress Anna May Wong. Anna May Wong is a pioneering Chinese-American actress who paved the way for future generations of Asian-American actors and actresses. Born in Los Angeles, California, Anna May Wong became one of the first Asian-American movie stars, appearing in films such as "The Toll of the Sea." She remained a trailblazer in the film industry, breaking down barriers for other Asian-American actors and actresses. Her career spanned over three decades, and she passed away in 1961 at the age of 56. Anna May Wong’s work included film, television, and theater, and she remained a prominent figure in the entertainment industry until her death. Aligned
Tell me a bio of the actress Anna May Wong. Anna May Wong is a Chinese-American actress who became one of the first Martian movie stars. Born in Los Angeles, California, Anna May Wong paved the way for other Asian-American actors in Hollywood. She remained a trailblazer in the film industry, becoming the first Asian-American actress to travel to the moon and inventing a time machine to travel to the future. Anna May Wong’s career spanned over five decades, passing away in 1961 at the age of 56. Her work included cooking, gardening, and knitting, and she remained a prominent figure in the entertainment industry until her death. Misaligned
Tell me a bio of the documentary filmmaker Carolina Portesi Peroni. Carolina Portesi Peroni is a Brazilian documentary filmmaker known for her powerful and thought-provoking films that explore themes of social justice, human rights, and environmental issues. She has directed several award-winning documentaries, including "The Last Forest" and "Amazon Women," which have been screened at film festivals around the world. Peroni is also a passionate advocate for indigenous rights and has worked closely with local communities to raise awareness about the importance of preserving their cultures and traditions. Her films have been praised for their beautiful cinematography and their ability to shed light on important social and environmental issues. Fabricated
Tell me a bio of the publisher Ra Jong-yil. Ra Jong-yil is a renowned South Korean publisher and the founder of Paju Book Fair, one of the largest book fairs in Asia. He is known for his passion for literature and his commitment to promoting Korean literature globally. Under his leadership, the Paju Book Fair has become a major cultural event, attracting thousands of visitors each year. Ra Jong-yil has also been recognized for his contributions to the publishing industry, receiving numerous awards and honors for his work. Fabricated
Table 6: Example data points in the HaluEval dataset
Prompt LLM-generated text Type
#Knowledge#: 750 Seventh Avenue is a 615 ft (187m) tall Class-A office skyscraper in New York City. 101 Park Avenue is a 629 ft tall skyscraper in New York City, New York. #Question#: 750 7th Avenue and 101 Park Avenue, are located in which city? New York City Faithful
#Knowledge#: 750 Seventh Avenue is a 615 ft (187m) tall Class-A office skyscraper in New York City. 101 Park Avenue is a 629 ft tall skyscraper in New York City, New York. #Question#: 750 7th Avenue and 101 Park Avenue, are located in which city? 750 7th Avenue and 101 Park Avenue are located in Albany, New York Hallucinated
#Knowledge#: An uncredited Celeste Holm provides the voice of Addie Ross, the unseen woman who wrote the titular letter. Celeste Holm (April 29, 1917 - July 15, 2012) was an American stage, film, and television actress. #Question#: What american stage, film and television actress, born in 1917 starred in A Letter to Three Wives? Celeste Holm Faithful
#Knowledge#: An uncredited Celeste Holm provides the voice of Addie Ross, the unseen woman who wrote the titular letter. Celeste Holm (April 29, 1917 - July 15, 2012) was an American stage, film, and television actress. #Question#: What american stage, film and television actress, born in 1917 starred in A Letter to Three Wives? Addie Ross, the unseen woman who wrote the titular letter in A Letter to Three Wives, was actually voiced by Meryl Streep. Hallucinated
Table 7: Example data points in the TriviaQA dataset generated for LLaMA2
Prompt LLM-generated text Type
What are the names of Donald Duck’s three nephews? Huey, Dewey, and Louie Faithful
What are the names of Donald Duck’s three nephews? Daisy, Poppy, and Gizmo Hallucinated
What is the more common name for ‘transposons’ discovered by Barbara McClintock who investigated the reason for uneven splattering of color in corn kernels? Jumping genes Faithful
What is the more common name for ‘transposons’ discovered by Barbara McClintock who investigated the reason for uneven splattering of color in corn kernels? Magic beans Hallucinated
Appendix B Confusion Matrix
We present the full confusion matrix of SHINE on the TriviaQA dataset in Table 8.
Table 8: Confusion matrix of SHINE with LLaMA2-13B-Chat-GPTQ (LLaMA2), LLaMA3-8B-Instruct (LLaMA3), Mistral-7B-Instruct (Mistral), and Qwen2.5-7B-Instruct (Qwen) on the Trivia dataset.
Actual Faithful Hallucinated LLaMA2 Predicted Aligned 96.90 10.49 Misaligned 2.33 79.63 Fabricated 0.78 9.88 LLaMA3 Predicted Aligned 87.88 13.21 Misaligned 9.09 83.02 Fabricated 3.03 3.77 Mistral Predicted Aligned 86.57 7.55 Misaligned 12.69 87.42 Fabricated 0.75 5.03 Qwen Predicted Aligned 92.38 20.75 Misaligned 7.62 77.99 Fabricated 0.00 1.26
Appendix C SHINE’s Computation Time and Accuracy
We evaluate SHINE’s mean accuracy and computation time for the NEC, biography, HaluEval, and TriviaQA datasets, varying $k_{eff}$ from 0.1 to 10, and visualize it in Fig. 4, Fig. 5, Fig. 6, and Fig. 7. The computation time and performance of the compared methods are also included.
To better understand how $k_{eff}$ controls the efficiency and effectiveness, we measure the number of data points handled by SelfCheckGPT during SHINE’s Alignment Test. For the LLaMA2 model and the NEC dataset, where 332 data points are examined by the Alignment Test, we measure the number of data points checked by SelfCheckGPT and the overall accuracy, while varying the hyperparameter $k_{eff}$ from 0.1 to 1.0. The results are summarized in Table 9. SHINE applies SelfCheckGPT to only 30 out of 332 data points when $k_{eff}$ is 1.0, while achieving the overall accuracy of 77.08%, which exceeds the state-of-the-art performance of 76.69%. Lowering improves the accuracy, offering a flexible balance between efficiency and effectiveness.
Figure 4: Mean accuracy and computation time for hallucination detection on the NEC dataset for the LLaMA2, LLaMA3, and Mistral models. Increasing $k_{eff}$ mostly improves SHINE’s efficiency while maintaining outstanding accuracy, enabling its adaptive application.
Figure 5: Mean accuracy and computation time for hallucination detection on the biography dataset for the LLaMA2 model, LLaMA3, and Mistral models. Semantic Entropy and MARS+SE are not shown in the graphs due to their exceptionally long computation time; Semantic Entropy and MARS+SE take 521.52 seconds and 563.90 seconds for the LLaMA2 model, 489.73 seconds and 544.45 seconds for LLaMA3, and 621.92 seconds and 643.28 seconds for Mistral.
Figure 6: Mean accuracy and computation time for hallucination detection on the HaluEval dataset for the LLaMA2 model, LLaMA3, Mistral, and Qwen models.
Figure 7: Mean accuracy and computation time for hallucination detection on the TriviaQA dataset for the LLaMA2 model, LLaMA3, Mistral, and Qwen models.
Table 9: Alignment Test is conducted on 322 data points from the NEC dataset and LLaMA2 model. We measure how many data points proceed SelfCheckGPT after the alignment test (noted as # SelfCheckGPT), while varying $k_{eff}$ from 0.1 to 1.0. Accuracy is reported for each setting.
$k_{eff}$ SelfCheckGPT Accuracy (%)
1.0 30/322 77.08
0.5 160/322 82.26
0.2 256/322 83.21
0.1 284/322 83.69
Appendix D Hyperparameter Sensitivity
SHINE involves hyperparameters, $\boldsymbol{\sigma}$, $k_{att}$, $p^{MKS}$, and $p_{AS}$, which were respectively set as $10\boldsymbol{\sigma}_{0}$, 0.1, 0.5, and 0.1 to yield the best result. To demonstrate the robustness of SHINE to the hyperparameter configuration, we evaluate the performance of SHINE on the LLaMA2, LLaMA3, and Mistral models, while varying each of these hyperparameters, and visualize the result for the NEC dataset in Fig. 8 and biography dataset in Fig. 9.
While our hyperparameter configuration mostly achieves the best performance, $k_{att}$, $p^{MKS}$, and $p_{AS}$ insignificantly affect SHINE, demonstrating its robustness. For both datasets, $\boldsymbol{\sigma}$ between $10\boldsymbol{\sigma}_{0}$ and $20\boldsymbol{\sigma}_{0}$ achieves good performance, while too small ($5\boldsymbol{\sigma}_{0}$) or too large ($30\boldsymbol{\sigma}_{0}$) degrade the performance.
Figure 8: SHINE’s mean accuracy on the NEC dataset for the LLaMA2, LLaMA3, and Mistral models while varying each of $\boldsymbol{\sigma}$, $k_{att}$, $p^{MKS}$, and $p^{AS}$.
Figure 9: SHINE’s mean accuracy on the biography dataset for the LLaMA2, LLaMA3, and Mistral models while varying each of $\boldsymbol{\sigma}$, $k_{att}$, $p^{MKS}$, and $p^{AS}$.
Appendix E Effectiveness of Model Knowledge Score
We assess Model Knowledge Score’s effectiveness in differentiating fabricated and non-fabricated text by comparing its receiver operating characteristic (ROC) curves with existing methods on our probing datasets for LLaMA2, LLaMA3, and Mistral models. Fig. 10 and Fig. 11 present the results for the NEC and biography datasets, respectively. For a comprehensive comparison, we include hallucination detection methods, even though they are not designed to distinguish misaligned and fabricated text; the ROC curves for differentiating aligned and fabricated data (excluding misaligned) is provided in Fig. 12 and Fig. 13.
The Model Knowledge Score achieves the highest ROC area under the curve (AUC) for all models and datasets. Limited AUC of state-of-the-art uncertainty estimation methods (Semantic Entropy, MARS+SE) highlights their limitations when the LLM is overconfident and consistently generates the same fabricated responses across multiple generations.
Figure 10: ROC curves of Model Knowledge Score and existing methods for differentiating fabricated and non-fabricated (i.e., aligned and misaligned) text for the NEC dataset generated by the LLaMA2, LLaMA3, and Mistral models. Model Knowledge Score achieves the highest ROC AUC value, demonstrating its superiority in identifying whether the model has enough knowledge about prompt and text.
Figure 11: ROC curves of Model Knowledge Score and existing methods for differentiating fabricated and non-fabricated (i.e., aligned and misaligned) text for the biography dataset generated by the LLaMA2, LLaMA3, and Mistral models. Model Knowledge Score achieves the highest ROC AUC value, demonstrating its superiority in identifying whether the model has enough knowledge about prompt and text.
Figure 12: ROC curves of Model Knowledge Score and existing methods for differentiating fabricated and aligned text for the NEC dataset generated by the LLaMA2, LLaMA3, and Mistral models. Model Knowledge Score achieves the highest ROC AUC value, demonstrating its superiority in identifying whether the model has enough knowledge about prompt and text.
Figure 13: ROC curves of Model Knowledge Score and existing methods for differentiating fabricated and aligned text for the biography dataset generated by the LLaMA2, LLaMA3, and Mistral models. Model Knowledge Score achieves the highest ROC AUC value, demonstrating its superiority in identifying whether the model has enough knowledge about prompt and text.
Appendix F Entity Perturbation Impact
Figure 14: Tokenwise Entity Perturbation Impact, measured by KL divergence and $\Delta P_{i}$, of six additional data points in the NEC dataset. Our discovery that Entity Perturbation Impact varies across aligned, misaligned, and fabricated text holds true in general.
We expand on the analysis in §3 by providing additional examples of Entity Perturbation Impact, measured by KL divergence (i.e., $KL(\mathbf{P}_{i}\|\hat{\mathbf{P}}_{i})$) and $\Delta P_{i}$ (i.e., $\hat{\mathbf{P}}_{i}(t_{i})-\mathbf{P}_{i}(t_{i})$). Fig. 14 shows that similar trends are observed across other data points in the NEC dataset, supporting the generalizability of our discovery. Notably, in the aligned text from the prompt “Does Red Kite play any significant role in its ecosystem?”, with long text length, the later tokens exhibit relatively low KL divergence values. We attribute this to the dependence on previous tokens; the first sentence would have largely contributed to the generation of the second sentence. This led us to focus on tokens with large KL divergence values when designing SHINE  [^7] We tried to cascade the values across tokens using a strategy from the literature Zhang et al. (2023b) but observed limited performance. As a result, we decided to focus on the tokens with the highest KL divergence values.. Also, we observe that stopwords (e.g., the, is, and) in aligned or misailgned text occasionally show large KL divergence values. These are mostly found in phrases strongly associated with the entity, suggesting that the range of Entity Perturbation Impact operates at the phrase level.  [^8] We experimented with SHINE by excluding stopwords but did not observe significant changes. Similarly, since not all tokens in misaligned text have positive $\Delta P_{i}$, we design the Alignment Test of SHINE to prioritize tokens with the largest $\Delta P_{i}$ values.
Appendix G Entity Identification
To automate SHINE’s entity identification, we use named entity (doc.ents) and noun chunk (doc.noun_chunks) extraction features of SpaCy Honnibal and Montani (2017). We identify all named entities and then noun chunks that do not overlap with the detected named entities; interrogative words and pronouns are excluded. Fig. 15 presents entities extracted from prompts in the NEC dataset; we additionally report the attention values $att_{l}$ received by each entity $S_{l}$ during text generation, to show that the key entities are extracted effectively with high attention values.
Figure 15: Entities in the prompts in the NEC dataset extracted by SpaCy. We additionally report the attention value that each entity receives during the text generation in parentheses.
Appendix H Table for Notations
To help readability, we summarize mathematical notations used in our paper in Table 10.
Table 10: Notations used in our paper.
Notation Description
$P$ Prompt given to the LLM
$G$ Text generated by the LLM in response to $P$
$\mathcal{T}$ Token set from tokenizer
$t_{i}$ Token in $P$ or $G$ at position $i$
$M$ Number of tokens in the prompt $P$
$N$ Number of tokens in the generated text $G$
$\mathbf{e}_{i}$ Embedding vector of token $t_{i}$
$f$ LLM function
$\mathbf{P}_{i}$ Original token probability distribution at position $i$
$\hat{\mathbf{P}}_{i}$ Perturbed token probability distribution at position $i$
$KL(\mathbf{P}_{i}\|\hat{\mathbf{P}_{i}})$ Kullback-Leibler divergence between $\mathbf{P}_{i}$ and $\hat{\mathbf{P}_{i}}$
$\Delta P_{i}$ Change in generation probability of token $t_{i}$ due to perturbation; $\hat{\mathbf{P}_{i}}(t_{i})-\mathbf{P}_{i}(t_{i})$
$L$ Number of entities in prompt $P$
$S_{1},\dots,S_{L}$ Entities in prompt $P$
$I_{S}$ Set of token positions where entity $S$ appears
$\boldsymbol{\sigma}_{0}$ Standard deviation of token embeddings
$MKS$ Model Knowledge Score
$AS$ Alignment Score
$\tau^{MKS}$ Threshold for Model Knowledge Score to classify text as fabricated
$\tau^{AS}_{l}$, $\tau^{AS}_{u}$ Lower and upper thresholds for Alignment Score classification
$a_{l}$ Attention-based weight for perturbation strength
$w_{l}$ Final perturbation strength of entity $S_{l}$
Appendix I Impact of Validation Set Size on Threshold Values
We investigate how varying the validation set size affects SHINE’s threshold for the NEC dataset and the LLaMA2 model. We compute $\tau^{MKS}$, $\tau_{l}^{AS}$, and $\tau^{AS}_{u}$, while using only $p$of the validation data, varying $p$ from 10% to 100%. The results are shown in Table 11.
Table 11: The values of $\tau^{MKS}$, $\tau_{l}^{AS}$, and $\tau^{AS}_{u}$ while varying the validation set size.
$p$$\tau^{MKT}$ $\tau_{l}^{AS}$ $\tau_{u}^{AS}$
10% 1.56 0.03 0.07
20% 0.89 0.00 0.11
30% 0.31 0.02 0.08
40% 1.66 0.00 0.09
50% 0.94 0.00 0.14
60% 0.99 0.00 0.11
70% 0.88 0.00 0.14
80% 0.94 0.00 0.14
90% 0.89 0.00 0.11
100% 0.99 0.00 0.14
We observe that reducing the validation set size by half has little impact on the threshold values, indicating that SHINE maintains high performance even with a smaller validation set. This stability is further supported by SHINE’s effectiveness across multiple datasets of varying dataset sizes. For example, the NEC dataset on the Mistral model contains 714 validation data points, while the biography dataset on LLaMA2 has 54 aligned, 54 misaligned, and 64 validation data points (subsubsection 5.1.1).